solve the system of equations by graphing then classify the system: y=-x-8 9x-8y=-38
Answer by mananth(16946)

1 x + 1 y = -8 .............1
9 x -8 y = -38 .............2

equation1
when x= 0 y= -8 ( 0 , -8 )
when y = 0 x= -8 ( -8 , 0 )
Equation 2
when x= 0 y= 4.75 ( 0 , 4.75 )
when y = 0 x= -4.22 ( -4.22 , 0 )


Solution ( -6,-2)